# 3]Sorbonne Université, CNRS, Université de Paris, Inria, Laboratoire Jacques-Louis Lions UMR 7598, 75005 Paris, France

### v1

@@ Abstract
Populations are heterogeneous, deviating in numerous ways. Phenotypic diversity refers to the range of traits or characteristics across a population, where for cells this could be the levels of signalling, movement and growth activity, etc. Clearly, the phenotypic distribution – and how this changes over time and space – could be a major determinant of population-level dynamics. For instance, across a cancerous population, variations in movement, growth, and ability to evade death may determine its growth trajectory and response to therapy. In this review, we discuss how classical partial differential equation (PDE) approaches for modelling cellular systems and collective cell migration can be extended to include phenotypic structuring. The resulting non-local models – which we refer to as phenotype-structured partial integro-differential equations (PS-PIDEs) – form a sophisticated class of models with rich dynamics. We set the scene through a brief history of structured population modelling, and then review the extension of several classic movement models – including the Fisher-KPP and Keller-Segel equations – into a PS-PIDE form. We proceed with a tutorial-style section on derivation, analysis, and simulation techniques. First, we show a method to formally derive these models from underlying agent-based models. Second, we recount travelling waves in PDE models of spatial spread dynamics and concentration phenomena in non-local PDE models of evolutionary dynamics, and combine the two to deduce phenotypic structuring across travelling waves in PS-PIDE models. Third, we discuss numerical methods to simulate PS-PIDEs, illustrating with a simple scheme based on the method of lines and noting the finer points of consideration. We conclude with a discussion of future modelling and mathematical challenges.
@@ 1 Introduction
Here the dependency on $(t,{\bm{x}})$ of ${\bm{A}_{\bm{x}}},D_{\bm{x}}$ and $F$ may also be mediated by $\rho$ itself, introducing non-linearities in the equation. The kinetics function $F$ describes proliferation and death, while cell migration is split into two components: an advective term with advective velocity $A_{\bm{x}}$, to describe directed movements, e.g. from guidance cues in the environment; a diffusive term with diffusion coefficient $D_{\bm{x}}$, to describe undirected (random) movements. The above can of course be extended to systems of coupled equations to define key additional dependent variables. A vast range of models have been built with equations that conform to this structure, including many classical systems found within standard textbooks, e.g. [191].
However, it has long been recognised that populations are rarely (if ever) homogeneous – some members may move faster, some may proliferate more quickly, etc – with potentially important consequences. Further, our technological capacity to finely discriminate the various forms of cellular heterogeneity is ever increasing: a repertoire of markers that highlight progression through the cell cycle, allowing an ‘age distribution’ to be tracked [84]; super resolution microscopy [136], that allows cell shape, size and structure to be recorded with unprecedented detail; protein and mRNA staining techniques (flow cytometry, immunocytochemistry, mass spectronomy, and so forth, see Figure 2(c)) provide information on the distribution of traits at a molecular level [216, 124]. In numerous systems, the sharpened picture afforded by these techniques has heightened our understanding into how population distributions are shaped by heterogeneity. To provide a couple of apposite examples: much recent attention has focused on the form of invading cancers, and their structuring from more migratory ‘leader’ cells at the front to more proliferative ‘follower’ cells at the rear [244], see Figure 2(a); microfabricated mazes reveal the phenotypic structure within E. coli bacteria populations, via sorting them according to their ability to respond to chemical gradients by chemotaxis [219], see Figure 2(b).
@@
We will review briefly this literature, but then turn our attention to mathematical methods and tools that can be used to model and analyse the evolving spatiotemporal distribution of a (cell) population, where the population members also span a (potentially broad) spectrum of phenotypes or traits: distinct traits could refer to differences in proliferative potential, their capacity to migrate, their ability to evade death, and so on. Phenotype variation within such models may range from a (relatively) simplistic binary description, to phenotypes extending across some high-dimensional continuum of states, see Figure 2(c). Models of this nature have become increasingly influential in recent years, a natural response to the large number of experimental studies that have revealed the role of different phenotypes when driving large scale population dynamics.
@@
The models we review here represent the extension of the basic reaction-advection-diffusion framework (1) to accommodate any impacts from phenotypic structuring. First, this involves including dependencies on the phenotypic state ${\bm{y}}$ on the various terms introduced above, so that the characteristic of a particular phenotype (e.g. higher proliferation, lower migration) can be incorporated. Second, it requires the incorporation of additional terms to describe potential transitions through phenotypic space. Specifically, we will review models built on the following structure:
@@
where the dependency on $(t,{\bm{x}},{\bm{y}})$ of ${\bm{A}_{\bm{x}}}$, $D_{\bm{x}}$, ${\bm{A}_{\bm{y}}}$, $D_{\bm{y}}$, and $F$ may be mediated by $\rho$ and $n$, introducing non-linearities and non-localities into the equation. Analogous to movements through space, phenotypic transitions can potentially occur in either a directed (e.g. factors in the environment that direct cells to switch from one phenotypic state to another) or undirected (e.g. a spontaneous mutation) fashion. Consequently, we introduce the two additional functions ${\bm{A}}_{\bm{y}}$ and $D_{\bm{y}}$ to describe directed and undirected transitions, respectively. Noting that the above equation extends the partial differential equation (PDE) (1) through both additional phenotype structuring and a potential dependence on the (total) density $\rho$, we refer to this as a phenotype-structured partial integro-differential equation (PS-PIDE).
The remainder of the review is laid out as follows. In the next section we provide a select historical overview, charting a few key modelling studies that have laid groundwork for the PS-PIDE models under review here. In Section 3 we review the extension of classic PDE models of cell motion of form (1) to the PS-PIDE form (2), focusing on some prototypical examples. In Section 4 we discuss the methods that allow these models to be derived from agent-based models and tools that enable their analysis, particularly in the context of travelling wave behaviour and concentration phenomena, and numerical simulation. Finally, we conclude with a brief discussion, organised into a series of future challenges for the field.
@@ 2 Structured population modelling
Models that incorporate population structuring have been around for more than a century and a broad range of approaches have been developed. Here we offer a short summary, concentrating on works most directly related to the PS-PIDE models described above (for a wider discussion, we refer to Hoppenstaedt 129, Perthame 202, Webb 249, Metz and Diekmann 181, Charlesworth 55, Cushing 68, Caswell 47, Kot 145, Auger et al. 12).
@@ 2.1 Structured populations and early modelling
Population counts date almost 6,000 years ago to early Mesopotamian civilisations and, subsequently, cultures from Egypt to China [123]. Surveys served purposes from economic (how much tax can we collect?) to militaristic (how many can fight?), but lacked information on population structure: usually, only the numbers of (able) men were recorded with women, children, or the infirm unstated [^3] Some records from China dating to the 5th century AD did indicate ages, sexes, and relationships of household members, see [82].. Censuses that include information on population structure form a (comparatively) recent undertaking, sometimes part-triggered by population growth concerns: the first (modern) UK census of 1801 took place in the aftermath of the controversy that surrounded Malthus’ 1798 essay on population growth [178]. Over time, censuses were expanded to collect [^4] UK Censuses in the latter half of the 19th century even requested numbers of ‘lunatics’, ‘imbeciles’, or ‘idiots’ living in a household, leading the Registrar General in 1881 to question the value of this information [196]: “It is against human nature to expect a mother to admit her young child to be an idiot, however much she may fear this to be true. To acknowledge the fact is to abandon all hope.”. greater information on population structure.
@@
Faced with this data, the question of how to describe a population is key to modelling. Early models – such as exponential growth ideas expounded by the likes of [178], or the logistic growth model of [239] – describe only total population over time, yet these models remain popular when structured information is missing, considered negligible, or dismissed. Models that include structuring date to the work of Alfred Lotka [^5] Well-known for the Lotka-Volterra equations, but demographics formed his main research area and later worked for the Metropolitan Life Insurance Company. The Lotka-Volterra model can also be regarded as a structured population model, structured into predators and prey. on demography in the early 1900s [170, 169, 221]. Lotka’s model took the form of a renewal equation and described a population’s evolving age distribution; analysis was used to determine whether a population would evolve to a stable age distribution.
@@ 2.2 Age-structured models
In (4)1, age and time progress in tandem and the right-hand side describes loss or death, with rate $\mu$, where $\rho(t)$ is the (total) population size at time $t$. The condition at $t=0$ is the initial age distribution of the population, i.e. some non-negative function on $[0,\infty)$. The condition at $a=0$ represents births, with $\beta$ the birth rate due to individuals of age $a$. For cell proliferation $\beta$ and $\mu$ may be linked: for example, $\beta=2\mu$ if a proliferation leads to two identical daughter cells – we refer to [202] for more discussion on applications of similar models to cell populations. The model (4) is of form (2), but with no space dependency and linear advection describing state transitioning. Dynamics predict the evolving age structure; for example, in Figure 2 a formulation is shown whereby early waves of distinct generations eventually give way to a stable distribution (but not necessarily stable population size).
@@
The different names given to (4) highlight its repeated discovery: beyond those of [180] and [240] [^8] Heinz von Foerster was a biophysicist with wide-ranging interests, including population growth; his “Doomsday equation” predicted that human population growth rate would become infinite (on November 13th, 2026) [241]. Tongue was firmly in cheek here, as the exact day coincided with what would have been his birthday. it was derived by [141] and [223]. Different derivations were slightly different in form – for example, a range from $\mu\equiv 0$ in [223] to a non-local dependency $\mu(\rho)$ in [114]. The derivation in [180] is the first documented, where age-structuring was used to illustrate [^9] Remarkably, the study of [180] used more than ten case applications, often based on data, in examples that included: house to house cancer incidence rates in Luckau, Germany; cholera epidemics in India; bacteria ingestion by leukocytes; infection and relapse in malaria; and, bacteria-antibody conflict, amongst several others. the compartmental approach when both time and the population state variable could be regarded as continuous. The state space then represents age (we change $c$ to $a$ in (3)), age and time are absolutely correlated, and ageing is (unfortunately) a one-way process: we can therefore set $l\equiv\tau$ and $F_{t}^{a\rightarrow a}=F_{t}^{a+\tau\rightarrow a}=0$. Not all individuals progress to the next age class (as some will die), leading to $F_{a}^{a-\tau\rightarrow a}=1-\mu\tau$ where $\mu\tau$ is the fraction of the population with age $a-\tau$ at time $t$ that die before they reach age $a$ at time $t+\tau$ Substituting these into (3), expanding about $(t,a)$ and taking the limit as $\tau\rightarrow 0$, one formally arrives at (4)1 under continuity assumptions.
@@
As noted, a derivation of (4) is also found in [141], for a stochastic birth-death process that accounted for age structuring of a population. These earlier derivations went somewhat under the radar (at least in the context of growth kinetics), with wider awareness not emerging until the 1960s – primarily to describe cellular growth – following the derivations of [223] and [240]. The naming after von Foerster appears to originate in [231, 230], and the 1960s witnessed its wider application to problems in cellular kinetics (e.g. Fredrickson and Tsuchiya 109, Trucco 231, 230, Rubinow 214) and population demographics (e.g. Hoppenstaedt 129, Keyfitz 142).
@@ 2.3 Age- and space-structured models
Age is one way a population can be structured, but model (4) can be straightforwardly extended to include other forms of heterogeneity. One approach would be to add a discrete structuring, such as coupling two equations of type (4)1 to account for female and male members of a population [104, 142]. Further continuous structure can be added, with early extensions to size and age structures proposed in [109, 21] and [225], for cell and animal populations, respectively. Generalisations to a $p$-dimensional “physiological state vector” were first made in [107], and we refer to [122, 249, 236, 250] for further discussion of high-dimensional structured population models. Other structuring variables may not obey the strict correlation of age with time, for example size may increase or decrease over time; as such, different transport terms may be needed to describe movements through the structuring variable state space.
Physical space is frequently one of the most crucial population structuring variables. Of course, mathematical models that account for spatiotemporal evolution have a long and illustrious history, frequently using the reaction-advection-diffusion form (1); we rarely refer to these as structured population models, though. Particularly pertinent here, however, are extensions of age-structured models to include spatial movement. These were first introduced by Gurtin and others [121, 115, 116], where a diffusive-type spatial movement was added to an age-structured population. Starting with model (4), but extending to a population $n(t,{\bm{x}},a)$ for time $t$, position ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ and age $a\in[0,\infty)$, yields
$\begin{cases}&\partial_{t}n+\partial_{a}n-\nabla_{\bm{x}}\cdot\left[D(a)\nabla n\right]=-\mu(t,{\bm{x}},a,n,\rho)n\,,\quad{\bm{x}}\in\mathcal{X}\ ,\;a\in(0,\infty)\ ,\\[5.0pt] &\displaystyle{\rho(t,{\bm{x}}):=\int_{0}^{\infty}n(t,{\bm{x}},a)\,{\rm d}a}\,,\\[5.0pt] &n(0,{\bm{x}},a)=n^{0}({\bm{x}},a)\,,\\[5.0pt] &\displaystyle{n(t,{\bm{x}},0)=\int_{0}^{\infty}\beta(t,{\bm{x}},a,n,\rho)n(t,{\bm{x}},a)\,{\rm d}a}\,,\\[5.0pt] &{\bm{\nu}}\cdot D(a)\nabla n=q(t,{\bm{x}},a)\mbox{ for }{\bm{x}}\in\partial\mathcal{X}\,,\end{cases}$
where ${\bm{\nu}}$ is the outer unit normal on the domain boundary $\partial\mathcal{X}$. The above is of PS-PIDE form, adapted and furnished with boundary and initial conditions relevant for a population structured in age and space. Structured models that include variation in space have subsequently received a huge amount of attention: a rather indiscriminate list includes [113, 248, 154, 103, 11, 13, 20, 19, 76, 80, 111, 81, 101, 83, 148], and involve applications from ecology to cancer growth. Indeed, the primary focus of the rest of this review will be on population models where the structuring is in space and phenotype. In this regard, we also note that variants of (4) for age- and phenotype-structured populations have recently been considered [195].

### v2

@@ Abstract
Populations are heterogeneous, deviating in numerous ways. Phenotypic diversity refers to the range of traits or characteristics across a population, where for cells this could be the levels of signalling, movement and growth activity, etc. Clearly, the phenotypic distribution – and how this changes over time and space – could be a major determinant of population-level dynamics. For instance, across a cancerous population, variations in movement, growth, and ability to evade death may determine its growth trajectory and response to therapy. In this review, we discuss how classical partial differential equation (PDE) approaches for modelling cellular systems and collective cell migration can be extended to include phenotypic structuring. The resulting non-local models – which we refer to as phenotype-structured partial differential equations (PS-PDEs) – form a sophisticated class of models with rich dynamics. We set the scene through a brief history of structured population modelling, and then review the extension of several classic movement models – including the Fisher-KPP and Keller-Segel equations – into a PS-PDE form. We proceed with a tutorial-style section on derivation, analysis, and simulation techniques. First, we show a method to formally derive these models from underlying agent-based models. Second, we recount travelling waves in PDE models of spatial spread dynamics and concentration phenomena in non-local PDE models of evolutionary dynamics, and combine the two to deduce phenotypic structuring across travelling waves in PS-PDE models. Third, we discuss numerical methods to simulate PS-PDEs, illustrating with a simple scheme based on the method of lines and noting the finer points of consideration. We conclude with a discussion of future modelling and mathematical challenges.
@@ 1 Introduction
Here the dependency on $(t,{\bm{x}})$ of ${\bm{A}_{\bm{x}}},D_{\bm{x}}$, and $F$ may also be mediated by $\rho$ itself, introducing non-linearities in the equation. The kinetics function $F$ describes proliferation and death, while cell migration is split into two components: an advective term with advective velocity $A_{\bm{x}}$, to describe directed movements, e.g. from guidance cues in the environment; a diffusive term with diffusion coefficient $D_{\bm{x}}$, to describe undirected (random) movements. The above can of course be extended to systems of coupled equations to define key additional dependent variables. A vast range of models have been built with equations that conform to this structure, including many classical systems found within standard textbooks, e.g. [195].
However, it has long been recognised that populations are rarely (if ever) homogeneous – some members may move faster, some may proliferate more quickly, etc – with potentially important consequences. Further, our technological capacity to finely discriminate the various forms of cellular heterogeneity is ever increasing: a repertoire of markers that highlight progression through the cell cycle, allowing an ‘age distribution’ to be tracked [85]; super resolution microscopy [139], that allows cell shape, size, and structure to be recorded with unprecedented detail; protein and mRNA staining techniques (flow cytometry, immunocytochemistry, mass spectronomy, and so forth, see Figure 2(c)) provide information on the distribution of traits at a molecular level [226, 134]. In numerous systems, the sharpened picture afforded by these techniques has heightened our understanding into how population distributions are shaped by heterogeneity. To provide a couple of apposite examples: much recent attention has focused on the form of invading cancers, and their structuring from more migratory ‘leader’ cells at the front to more proliferative ‘follower’ cells at the rear [243], see Figure 2(a); microfabricated mazes reveal the phenotypic structure within E. coli bacteria populations, via sorting them according to their ability to respond to chemical gradients by chemotaxis [218], see Figure 2(b).
@@
We will review briefly this literature, but then turn our attention to mathematical methods and tools that can be used to model and analyse the evolving spatiotemporal distribution of a (cell) population, where the population members also span a (potentially broad) spectrum of phenotypes or traits: distinct traits could refer to differences in proliferative potential, capacity to migrate, ability to evade death, and so on. Phenotype variation within such models may range from a (relatively) simplistic binary description, to phenotypes extending across some high-dimensional continuum of states, see Figure 2(c). Models of this nature have become increasingly influential in recent years, a natural response to the large number of experimental studies that have revealed the role of different phenotypes when driving large scale population dynamics.
@@
The models we review here represent the extension of the basic reaction-advection-diffusion framework (1) to accommodate any impacts from phenotypic structuring. First, this involves including dependencies on the phenotypic state ${\bm{y}}$ in the various terms introduced above, so that the characteristic of a particular phenotype (e.g. higher proliferation, lower migration) can be incorporated. Second, it requires the incorporation of additional terms to describe potential transitions through phenotypic space. Specifically, we will review models built on the following structure:
@@
where the dependency on $(t,{\bm{x}},{\bm{y}})$ of ${\bm{A}_{\bm{x}}}$, $D_{\bm{x}}$, ${\bm{A}_{\bm{y}}}$, $D_{\bm{y}}$, and $F$ may be mediated by $\rho$ and $n$, introducing non-linearities and non-localities into the equation. Analogous to movements through space, phenotypic transitions can potentially occur in either a directed (e.g. factors in the environment that direct cells to switch from one phenotypic state to another) or undirected (e.g. a spontaneous mutation) fashion. Consequently, we introduce the two additional functions ${\bm{A}}_{\bm{y}}$ and $D_{\bm{y}}$ to describe directed and undirected transitions, respectively. Noting that the above equation extends the partial differential equation (PDE) (1) through both additional phenotype structuring and a potential dependence on the (total) density $\rho$, we refer to this as a phenotype-structured partial differential equation (PS-PDE).
Reading guide
This article covers both review and tutorial elements and different components may be of varying interest to different readers. We provide the following signpost.
  • Section 2 offers an overview of structured population models. This serves to place the model (2) in a broader and historical context, but readers not interested in this aspect can skip this section.
  • Section 3 covers three case studies, aimed to show how well-known PDE models for cell movement can be extended into a PS-PDE form. Specifically, we show this for a Fisher-KPP equation, a pressure-based movement model, and a taxis-based model.
  • Section 4 forms a tutorial, showing how standard tools and methods used to derive and analyse PDE models can be extended to their phenotype-structured counterparts. Specifically, we show a derivation of these continuum models from an on lattice agent-based model, illustrate how classical analyses of concentration phenomena and travelling waves can be extended to phenotype-structured models, and provide a straightforward scheme for solving such models. We note that such a scheme has been made publicly available on a repository.
  • Section 5 provides our perspectives on future challenges and directions for this field, covering elements that include applications, modelling, analysis, numerical simulation, and connecting to data.
@@ 2 Structured population modelling
Models that incorporate population structuring have been around for more than a century and a broad range of approaches have been developed. Here we offer a short summary, concentrating on works most directly related to the PS-PDE models described above (for a wider discussion, we refer to Hoppenstaedt 132, Perthame 205, Webb 252, Metz and Diekmann 188, Charlesworth 55, Cushing 69, Caswell 49, Kot 154, Auger et al 18).
@@ 2.1 Structured populations and early modelling
Population counts date almost 6,000 years ago to early Mesopotamian civilisations and, subsequently, cultures from Egypt to China [125]. Surveys served purposes from economic (how much tax can we collect?) to militaristic (how many can fight?), but lacked information on population structure: usually, only the numbers of (able) men were recorded with women, children, or the infirm unstated [^3] Some records from China dating to the 5th century AD did indicate ages, sexes, and relationships of household members, see [83].. Censuses that include information on population structure form a (comparatively) recent undertaking, sometimes part-triggered by population growth concerns: the first (modern) UK census of 1801 took place in the aftermath of the controversy that surrounded Malthus’ 1798 essay on population growth [183]. Over time, censuses were expanded to collect [^4] UK Censuses in the latter half of the 19th century even requested numbers of ‘lunatics’, ‘imbeciles’, or ‘idiots’ living in a household, leading the Registrar General in 1881 to question the value of this information [199]: “It is against human nature to expect a mother to admit her young child to be an idiot, however much she may fear this to be true. To acknowledge the fact is to abandon all hope.”. more information on population structure.
@@
Faced with these data, the question of how to describe a population is key to modelling. Early models – such as exponential growth ideas expounded by the likes of [183], or the logistic growth model of [242] – describe only total population over time, yet these models remain popular when structured information is missing, considered negligible, or dismissed. Models that include structuring date to the work of Alfred Lotka [^5] Well-known for the Lotka-Volterra equations, but demographics formed his main research area and later worked for the Metropolitan Life Insurance Company. The Lotka-Volterra model can also be regarded as a structured population model, structured into predators and prey. on demography in the early 1900s [179, 178, 222]. Lotka’s model took the form of a renewal equation and described a population’s evolving age distribution; analysis was used to determine whether a population would evolve to a stable age distribution.
@@ 2.2 Age-structured models
In [^8] In this manuscript we use the notation (m)i to refer to the equation on line $i$ for the group with number reference (m). (4)1, age and time progress in tandem and the right-hand side describes loss or death, with rate $\mu$, where $\rho(t)$ is the (total) population size at time $t$. The condition at $t=0$ is the initial age distribution of the population, i.e. some non-negative function on $[0,\infty)$. The condition at $a=0$ represents births, with $\beta$ the birth rate due to individuals of age $a$. For cell proliferation $\beta$ and $\mu$ may be linked: for example, $\beta=2\mu$ if a proliferation leads to two identical daughter cells – we refer to [205] for more discussion on applications of similar models to cell populations. The model (4) is of form (2), but with no space dependency and linear advection describing state transitioning. Dynamics predict the evolving age structure; for example, in Figure 2 a formulation is shown whereby early waves of distinct generations eventually give way to a stable distribution (but not necessarily stable population size).
@@
The different names given to (4) highlight its repeated discovery: beyond those of [185] and [102] [^9] Heinz von Foerster was a biophysicist with wide-ranging interests, including population growth; his “Doomsday equation” predicted that human population growth rate would become infinite (on November 13th, 2026) [247]. Tongue was firmly in cheek here, as the exact day coincided with what would have been his birthday. it was derived by [147] and [219]. Different derivations were slightly different in form – for example, a range from $\mu\equiv 0$ in [219] to a non-local dependency $\mu(\rho)$ in [121]. The derivation in [185] is the first documented, where age-structuring was used to illustrate [^10] Remarkably, the study of [185] used more than ten case applications, often based on data, in examples that included: house to house cancer incidence rates in Luckau, Germany; cholera epidemics in India; bacteria ingestion by leukocytes; infection and relapse in malaria; and, bacteria-antibody conflict, amongst several others. the compartmental approach when both time and the population state variable could be regarded as continuous. The state space then represents age (we change $c$ to $a$ in (3)), age and time are absolutely correlated, and ageing is (unfortunately) a one-way process: we can therefore set $l\equiv\tau$ and $F_{t}^{a\rightarrow a}=F_{t}^{a+\tau\rightarrow a}=0$. Not all individuals progress to the next age class (as some will die), leading to $F_{a}^{a-\tau\rightarrow a}=1-\mu\tau$ where $\mu\tau$ is the fraction of the population with age $a-\tau$ at time $t$ that die before they reach age $a$ at time $t+\tau$ Substituting these into (3), expanding about $(t,a)$, and taking the limit as $\tau\rightarrow 0$, one formally arrives at (4)1 under continuity assumptions.
@@
As noted, a derivation of (4) is also found in [147], for a stochastic birth-death process that accounted for age structuring of a population. These earlier derivations went somewhat under the radar (at least in the context of growth kinetics), with wider awareness not emerging until the 1960s – primarily to describe cellular growth – following the derivations of [219] and [102]. The naming after von Foerster appears to originate in [234, 235], and the 1960s witnessed its wider application to problems in cellular kinetics (e.g. Fredrickson and Tsuchiya 106, Trucco 234, 235, Rubinow 217) and population demographics (e.g. Hoppenstaedt 132, Keyfitz 149).
@@ 2.3 Age- and space-structured models
Age is one way a population can be structured, but model (4) can be straightforwardly extended to include other forms of heterogeneity. One approach would be to add a discrete structuring, such as coupling two equations of type (4)1 to account for female and male members of a population [105, 149]. Further continuous structure can be added, with early extensions to size and age structures proposed by [106, 24], and [225]. Generalisations to a $p$-dimensional “physiological state vector” were first made in [107], and we refer to [124, 252, 237, 251] for further discussion of high-dimensional structured population models. Other structuring variables may not obey the strict correlation of age with time, for example size may increase or decrease over time; as such, different transport terms may be needed to describe movements through the structuring variable state space.
Physical space is frequently one of the most crucial population structuring variables. Of course, mathematical models that account for spatiotemporal evolution have a long and illustrious history, frequently using the reaction-advection-diffusion form (1); we rarely refer to these as structured population models, though. Particularly pertinent here, however, are extensions of age-structured models to include spatial movement. These were first introduced by Gurtin and others [120, 122, 123], who added a diffusive-type spatial movement to an age-structured population. Starting with model (4), but extending to a population $n(t,{\bm{x}},a)$ for time $t$, position ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ and age $a\in[0,\infty)$, yields
$\begin{cases}&\partial_{t}n+\partial_{a}n-\nabla_{\bm{x}}\cdot\left[D(a)\nabla n\right]=-\mu(t,{\bm{x}},a,n,\rho)n\,,\quad{\bm{x}}\in\mathcal{X}\ ,\;a\in(0,\infty)\ ,\\[5.0pt] &\displaystyle{\rho(t,{\bm{x}}):=\int_{0}^{\infty}n(t,{\bm{x}},a)\,{\rm d}a}\,,\\[5.0pt] &n(0,{\bm{x}},a)=n^{0}({\bm{x}},a)\,,\\[5.0pt] &\displaystyle{n(t,{\bm{x}},0)=\int_{0}^{\infty}\beta(t,{\bm{x}},a,n,\rho)n(t,{\bm{x}},a)\,{\rm d}a}\,,\quad{\bm{x}}\in\mathcal{X}\,,\\[5.0pt] &{\bm{\nu}}\cdot D(a)\nabla n=q(t,{\bm{x}},a),\quad{\bm{x}}\in\partial\mathcal{X}\ ,\;a\in(0,\infty),\end{cases}$
where ${\bm{\nu}}$ is the outer unit normal on the domain boundary $\partial\mathcal{X}$. The above is of PS-PDE form, adapted and furnished with boundary and initial conditions relevant for a population structured in age and space. Structured models that include variation in space have subsequently received a huge amount of attention: a rather indiscriminate list includes [113, 250, 157, 99, 1, 2, 19, 20, 75, 84, 112, 79, 100, 76, 142], and involve applications from ecology to cancer growth. Indeed, the primary focus of the rest of this review will be on population models where the structuring is in space and phenotype. In this regard, we also note that variants of (4) for age- and phenotype-structured populations have recently been considered [198].